[Brahms-l] Shift report 20040207 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Sat Feb 07 2004 - 00:01:22 EST
------------------------------------------------------------
Date: Sat Feb  7 00:01:22 2004, Shift: 16:00-24:00
Supervisor: Ionut Arsene, with: 
Log book#15 pages 49 - 49

Shift Summary:
------------------------------------------------------------
 No beam for the entire shift. Quench Link Interlock and 
Tandem Spark problems.

We had also problems with some of our detectors (see E-log). 
The big tubes from BBC tripped.
Problems also with D2 because of overheating. D2 was put
back on, but the monitoring page for this magnet is not give 
readout values corectly.
 
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Feb 7 00:01:28 2004

This archive was generated by hypermail 2.1.8 : Sat Feb 07 2004 - 00:01:49 EST